You are evaluating two different cookie-baking ovens. The Pillsbury 707 costs $69,500, has a 5-year life, and has an annual OCF (after-tax) of-$11,200 per year. The Keebler CookieMunster costs $96,000, has a 7-year life, and has an annual OCF (after-tax) of-$9,200 per year If your discount rate is 12 percent, what is each machines EAC? (Negative amounts should be indicated by a minus sign. Round your answers to 2 decimal places.) EAC $ 23,860.77 Pillsbury 707 Keebler CookieMunster24,800.97
Answer #1

Equivalent Annual Costs (EAC) = Asset Price * [Discount Rate / 1-(1+ Discount Rate)^-n] + Annual Maintenance Charges.

EAC for Pillsbury 707 = {69500 * [0.12 / 1- (1+0.12)^-5] +11200 }= { 69500 * [ 0.12 / 1-0.5674]+11200} = {69500 * [0.12/0.4325] +11200}

= (69500 * 0.2774) +11200

= $ 30479.97.

EAC for Keebler CookieMunster = {96000 * [0.12 / 1- (1+0.12)^-7] +9200 }= { 96000 * [ 0.12 / 1-0.4523]+9200} = {96000 * [0.12/0.5476] +9200}

= (96000 * 0.2191) +9200

= $ 30235.30